OpenStudy (anonymous):

what polynomial has roots -2, -3i, and 3i?

OpenStudy (anonymous):

a polynomial with (x+ 2)(x^2 + 9) Can you multiply these together?

OpenStudy (anonymous):

x^3 + 2x^2 + 9x +18 satisfies the conditions

OpenStudy (anonymous):

\[\frac{\pi^2}{9e}x^3 + \frac{2\pi^2}{9e}x^2 + \frac{\pi^2}{e}x + \frac{2\pi^2}{e}\] appears to be a polynomial with those roots

OpenStudy (anonymous):

2x^3 + 4x^2 + 18x +36

OpenStudy (anonymous):

2ex^3 + 4ex^2 + 18ex +36e also satisfies the condition